Determine the amount of chloride in 10.0 g of barium chloride. 


1
Expert's answer
2021-01-11T03:50:04-0500

Solution:


M("BaCl_2") - 208 "\frac{g} {mole}"



In the given chemical compaunds, we have 2 chlorine atoms :


208 "\frac{g} {mole}" - 71 "\frac{g} {mole}"

10 g - x g


x= "\dfrac{10\cdot 71} {208} = 3.41 g"


Answer:


The amount of chlorine in the 10 g of barium chloride is 3.41 g.